# A Review of Bayesian Uncertainty Quantification in Deep Probabilistic Image Segmentation

### v5

@@ 6.1 Observer variability
To contextualize this phenomenon within the framework of uncertainty quantification, annotators can be regarded as models themselves. For example, consider $K$ separate annotators modeled through parameters $\bm{\phi}_{k}$ with $k=1,2,...,K$. For a simple segmentation task, it can be expected that $\text{Var}[\,p(\bm{\phi}_{k})\,]\rightarrow 0$. In other words, each annotator is consistent in his delineation and the intra-observer variability is low. For cases with consensus across experts, i.e. yielding negligible inter-observer variability, the marginal approaches to $\text{Var}[\,p(\bm{\phi})\,]\rightarrow 0$. Asserting these two assumptions, it is valid to simply consider a point estimate of the posterior. Yet, this is rarely the case in many real-life applications and, as such, explicitly modeling the involved distributions becomes necessary.
@@
$\text{GED}(P_{\mathbf{Y}},P_{\hat{\mathbf{Y}}})$ $=\mathbb{E}_{\mathbf{y},\mathbf{y}^{\prime}\sim P_{\mathbf{Y}}}[k(\mathbf{y},\mathbf{y}^{\prime})]+\mathbb{E}_{\hat{\mathbf{y}},\hat{\mathbf{y}}^{\prime}\sim P_{\hat{\mathbf{Y}}}}[k(\hat{\mathbf{y}},\hat{\mathbf{y}}^{\prime})]-2\cdot\mathbb{E}_{\mathbf{y}\sim P_{\mathbf{Y}}}\mathbb{E}_{\hat{\mathbf{y}}\sim P_{\hat{\mathbf{Y}}}}[k(\mathbf{y},\hat{\mathbf{y}})],$ (36)
@@
The most straightforward approach is to directly model the empirical stochasticity in the annotations with models that target the aleatoric uncertainty. Given that annotators have an intrinsically associated expertise, it can also be possible to modeling the annotator distribution. However, relying on the model to infer ambiguity in the parameters by observing the data, can become quite burdensome. This observation is reflected in literature, where we see that most approaches use conditional generative models to encapsulate observer variability (See Table 1). While conditional generative models have been successfully applied for the task at hand, it has been shown that such models, without explicit conditioning, do not encapsulate more subtle variations such as distinct labeling styles of annotators Zepf et al. (2023a). VAE-based Probabilistic U-Net is the most popular approached followed by its hierarchical variants (HPU-Net and PHI-Seg) and SSNs. More recently, the growing popularity of DDPMs is apparent in the field.
@@
VAE Gao et al. (2022); Kohl et al. (2019b); Valiuddin et al. (2021); Selvan et al. (2020); Bhat et al. (2023; 2022a); Valiuddin et al. (2024b); Qiu & Lui (2020); Valiuddin et al. (2024a); Viviers et al. (2023c); Schmidt et al. (2023); Long et al. (2021b); Zepf et al. (2023a; 2024); Hu et al. (2022); Viviers et al. (2023b); Savadikar et al. (2021); Philps et al. (2024)
HVAE Zhang et al. (2022b); Rafael-Palou et al. (2021); Kohl et al. (2019a); Baumgartner et al. (2019); Gantenbein et al. (2020)
@@ 6.2 Active Learning
The field of active learning (Settles, 2009; Ren et al., 2021) aims to reduce the costly annotation procedure by careful selection of the most informative samples for training (Settles, 2009). In this way, the required labeled data is minimized and training convergence can be accelerated due to reduced redundancy. Active Learning is an extensively researched topic in both traditional (Cohn et al., 1996) and deep Machine Learning (Gal et al., 2017b), including uncertainty-based approaches applied to image-segmentation pairs. In such methods, it is assumed that predictions with high uncertainty provide the most informative samples for training.
@@
Selecting the most informative samples for training contains two elements. Firstly, the uncertainty needs to be quantified. This is usually the predictive entropy Kasarla et al. (2019); Shen et al. (2021); Xie et al. (2022) or variance Ozdemir et al. (2021), but in some instances the epistemic uncertainty is explicitly estimated through Bayesian Active Learning by Disagreement (BALD) (Houlsby et al., 2011; Ma et al., 2024b; Shen et al., 2021). Simply stated, the reduction in posterior entropy indicates the informativeness of a data sample. In turn, this can also be formulated by the reduction in predictive entropy in the dual formulation. We can formally denote this as
@@
Note the similarity between the mutual information $I$ obtained from the predictive entropy decomposition in Equation (3), where both are the objective of epistemic uncertainty modeling. Secondly, to evaluate whether the active learning strategy is beneficial, the performance is mostly measured subject to increasingly stringent budget requirements (i.e. a percentage of the original dataset) Yang et al. (2017); Sourati et al. (2018); Kasarla et al. (2019); Mahapatra et al. (2018); Sinha et al. (2019); Siddiqui et al. (2020); Li & Alstrøm (2020); Kim et al. (2021); Shen et al. (2021); Burmeister et al. (2022); Ma et al. (2024b). Kahl et al. (2024) argue that Active Learning should be evaluated on image-level as humans dont annotate on pxiel-level. Assuming a saturated performance on in-distribution data (training cycle $t_{1}$), the author propose to measure the improved relative performance, $C$, subject to a second training sample ($t_{2}$) from selected samples
@@
Table 2: Methods used for the application of Active Learning
@@
MC dropout Kahl et al. (2024); Gorriz et al. (2017); Ozdemir et al. (2021); Hiasa et al. (2019); Li & Alstrøm (2020); Shen et al. (2021); Ma et al. (2024b); Gaillochet et al. (2023); Siddiqui et al. (2020); Sadafi et al. (2019)
@@
Since the nature of this problem involves identifying and reducing model ignorance, the quantification of epistemic uncertainty is most appropriate. Hence, as can be seen in Table 2, most of the literature use MC dropout of ensembles. Furthermore, using the plain SoftMax activation has also been commonly used. Often, solely relying on the most uncertain samples can result in samples with limited diversity and therefore samples selected on the representativeness Ozdemir et al. (2021); Shen et al. (2021); Wu et al. (2021); Mahapatra et al. (2018); Sinha et al. (2019). Another crucial modeling element pertains the aggregation of the pixel-level uncertainties. Therefore, closeness to boundary Gorriz et al. (2017); Kasarla et al. (2019); Ma et al. (2024b) or specific regions Gaillochet et al. (2023); Wu et al. (2021); Kasarla et al. (2019) are additionally incorporated when combining the uncertainty values across pixels. Siddiqui et al. (2020) determine the entropy through estimating the entropy across viewpoints in multi-view datasets. Gaillochet et al. (2023) argue the superiority of selecting samples based on batch-level uncertainty.
@@ 6.3 Model introspection
Erroneous predictions should correlate with the uncertainty in cases the source is assumed to be epistemic. In turn, this implies that the provided uncertainty can be used to to gauge prediction quality of a model. High uncertainty predictions can be discarded to ensure a level of model reliability. Hence, the described problem can also be considered as Out-of-Distribution (OOD) detection (Lambert et al., 2022; Holder & Shafique, 2021). Including information on localized uncertainty to the training objective has shown to improve generalization capabilities (Ozdemir et al., 2017; Bian et al., 2020; Iwamoto et al., 2021; Li et al., 2021). Note that using uncertainty to guide model training is closely related to Active Learning, which has been discussed in Section 6.2.
Table 3: Methods used for the application of Model Introspection
@@
TTA Kahl et al. (2024); Wang et al. (2019b); Whitbread & Jenkinson (2022); Roshanzamir et al. (2023); Dahal et al. (2020)
Ensembling Jungo et al. (2020); Mehrtash et al. (2020); Czolbe et al. (2021); Holder & Shafique (2021); Jungo & Reyes (2019); Ng et al. (2022); Hann et al. (2021); Jungo & Reyes (2019); Linmans et al. (2020); Pavlitskaya et al. (2020); Valada et al. (2017)
MC Dropout Jungo et al. (2020); Kahl et al. (2024); Whitbread & Jenkinson (2022); Kendall et al. (2016); Kampffmeyer et al. (2016); Dechesne et al. (2021); Morrison et al. (2019); Qi et al. (2023); Eaton-Rosen et al. (2018); Jungo et al. (2018a; b); Roy et al. (2018a; b); Mehrtash et al. (2020); Nair et al. (2020); Sander et al. (2019); Hasan & Linte (2022b); Camarasa et al. (2021); Roshanzamir et al. (2023); Sedai et al. (2018); Seeböck et al. (2019); DeVries & Taylor (2018); Czolbe et al. (2021); Hoebel et al. (2019); Bhat et al. (2021); Dahal et al. (2020); Antico et al. (2020); Lambert et al. (2022); Hasan & Linte (2022a); Jungo & Reyes (2019); Ng et al. (2022); Iwamoto et al. (2021); Hoebel et al. (2020)
@@
VAE Chotzoglou & Kainz (2019); Viviers et al. (2023a); Czolbe et al. (2021); Bian et al. (2020)
@@
The relationship between uncertainty and model accuracy has been formalized by Mukhoti & Gal (2018) through the Patch Accuracy vs Patch Uncertainty (PAvPU) metric. Firstly, the accuracy given a certain prediction, $p(\text{A}|\text{C})$, and secondly, the uncertainty given an inaccurate prediction $p(\text{U}|\text{I})$. Given a threshold $u_{T}$ that distinguishes certain from uncertain pixels or patches, we can define pixels that are accurate and certain, accurate and uncertain, inaccurate and certain, inaccurate and uncertain, denoted by $u_{\text{ac}}$, $u_{\text{au}}$, $u_{\text{ic}}$, $u_{\text{iu}}$, respectively. Consequently, the authors combine $p(\text{A}|\text{C})=n_{\text{ac}}/(n_{\text{ac}}+n_{\text{au}})$ and $p(\text{U}|\text{I})=n_{\text{ic}}/(n_{\text{ic}}+n_{\text{iu}})$ to obtain the Patch Accuracy vs Patch Uncertainty (PAvPU) metric, defined as
@@
It can be noted that uncertainty is usually only obtained on pixel basis, while crucial information can be present in structural statistics. Hence, Roy et al. (2018b) propose to use the Coefficient of Variation (CoV) addresses this by measuring structural uncertainty through dividing the volume variance over the mean for all samples. The authors also evaluate structural uncertainties by assuming predictions with the pair-wise average overlap between all respective samples.
@@
Considering Table 3, it can be observed that epistemic methods such as MC dropout and Ensembles are by far the most popular choice of uncertainty encapsulation method. However, Concrete dropout (Rumberger et al., 2020b; Mukhoti & Gal, 2018), M-heads (auxiliary networks) (Jungo & Reyes, 2019; Linmans et al., 2020), MoE (Pavlitskaya et al., 2020) Laplace Approximation (Zepf et al., 2023b), Variational Inference in 3D (LaBonte et al., 2019), and test-time augmentation (Dahal et al., 2020) have also been used for this application.
@@ 6.4 Model generalization
As mentioned earlier in Section 5.1, sampling new parameter permutations often improves segmentation performance, due to the model combining effect. In contrast to other downstream tasks, evaluation does not require specialized metrics besides quantifying the relative change in model performance. For instance, dropout layers at the deeper decomposition levels of the SegNet (Badrinarayanan et al., 2015) improves model performance (Kendall et al., 2016). In Table 4, we can see improved performance with MC Dropout and also in specific with Concrete Dropout (Mukhoti & Gal, 2018; Rumberger et al., 2020b). The improved generalization from ensembling has also shown to produce more calibrated outputs (Mehrtash et al., 2020). Ng et al. (2022) found that ensembling results in the best performance improvement, while BBB is more robust to noise distortions. In other work, orthogonality within and across convolutional filters of the ensemble is enforced through minimizing their cosine similarity, which reaped similar merits (Larrazabal et al., 2021b). Ensembling has also been performed using M-Heads (Hu et al., 2023; Jungo & Reyes, 2019). Nonetheless, individual models in conventional ensembles receive data in an unstructured manner. Hence, specific subsets of the data can also be assigned to particular models with MoEs (Pavlitskaya et al., 2020; Ji et al., 2021; Gao et al., 2022). Generative models, though typically used for other applications, have also led to performance improvements across a wide range of studies Saha et al. (2021a); Wolleb et al. (2021); Wu et al. (2023b); Amit et al. (2021); Bogensperger et al. (2023); Zbinden et al. (2023); Viviers et al. (2023a); Zepf et al. (2023a); Hu et al. (2022).
@@
MC dropout Mukhoti & Gal (2018); Rumberger et al. (2020b); Saha et al. (2021a); Zhang et al. (2022a); Whitbread & Jenkinson (2022); Huang et al. (2018); Mukhoti et al. (2020); Ng et al. (2022); Wickstrøm et al. (2020)
Ensembling Ng et al. (2022); Larrazabal et al. (2021b); Kamnitsas et al. (2018); Saha et al. (2021b); Hu et al. (2023); Ji et al. (2021)
TTA Wang et al. (2019a); Rakic et al. (2024); Whitbread & Jenkinson (2022)

### v8

@@ 6.1 Observer variability
For example, in critical applications like radiation therapy planning, the precise delineation of tumor contours is essential. However, inherent ambiguity in the image acquisition process often leads to significant inter-observer variability. Multiple expert radiologists might annotate the same tumor with different, yet equally plausible, boundaries. In this context, no single annotation can be considered definitively correct, and this irreducible uncertainty in the ground truth must be accounted for to mitigate potentially serious impacts on patient care. Similarly, the perception of an autonomous driving system can suffer from label-level uncertainty. Consider an object partially obscured at a distance. The sensor data may be insufficient to distinguish whether it is a pedestrian about to cross the road or a stationary object like a mailbox due to a genuine ambiguity in the scene itself. The decision-making process must handle this uncertainty, as the consequences of misclassifying a vulnerable road user are severe.
@@
To contextualize this phenomenon within the framework of uncertainty quantification, annotators can be regarded as models themselves. For example, consider $K$ separate annotators modeled through parameters $\bm{\phi}_{k}$ with k=1,2,,Kk=1,2,...,K. For a simple segmentation task, it can be expected that $\text{Var}[\,p(\bm{\phi}_{k})\,]\rightarrow 0$. In other words, each annotator is consistent in their delineation and the intra-observer variability is low. For cases with consensus across experts, i.e. yielding negligible inter-observer variability, the marginal approaches to $\text{Var}[\,p(\bm{\phi})\,]\rightarrow 0$. Asserting these two assumptions, it is valid to simply consider a point estimate of the posterior. Yet, this is rarely the case in many real-life applications and, as such, explicitly modeling the involved distributions becomes necessary.
@@
with marginals $P_{\*Y}$ and $P_{\hat{\*Y}}$, representing the true and predictive segmentation distribution, and a distance represented by kernel $k:\mathcal{Y}\times\mathcal{Y}\rightarrow\mathbb{R}$, usually the (1-IoU) or (1-Dice) score (squared by the GED). However, solely relying on the GED has been criticized due to undesirable inductive biases (Kohl et al., 2019a; Zepf et al., 2024). Hence, Kohl et al. (2019a) introduce an alternative metric known as Hungarian Matching (HM), which compares the predictions against the ground-truth labels through a cost matrix. This can be also formally denoted as finding the permutation matrix P\*P, subject to the objective
@@
where the elements of M\*M are $M_{i,j}=k(\mathcal{Y}_{i},\mathcal{\hat{Y}}_{j})$, and $N^{2}$ represents the number of elements within the matrix. Subsequently, the unique optimal coupling between the two sets that minimize the average cost is determined through a combinatorial optimization algorithm. Hu et al. (2022) use the Normalized Cross Correlation (NCC) to measure the similarity between the true and predictive set, which can formally be defined as
@@
The most straightforward approach is to directly model the empirical stochasticity in the annotations with stochasticity in the features. Given that annotators have an intrinsically associated expertise, it can also be possible to model the annotator distribution in the parameters. However, relying on the model to infer ambiguity in the parameters by observing the data, can become quite burdensome. This observation is reflected in literature, where we see that most approaches use conditional generative models to encapsulate observer variability (See Table 1). While conditional generative models have been successfully applied for the task at hand, it has been shown that such models, without explicit conditioning, do not encapsulate more subtle variations such as distinct labeling styles of annotators (Zepf et al., 2023a). VAE-based Probabilistic U-Net is the most popular approached followed by its hierarchical variants (HPU-Net and PHI-Seg) and SSNs. More recently, the growing popularity of DDPMs is apparent in the field.
@@
VAE Gao et al. (2022); Kohl et al. (2019b); Valiuddin et al. (2021); Selvan et al. (2020); Bhat et al. (2023; 2022a); Valiuddin et al. (2024b); Qiu and Lui (2020); Valiuddin et al. (2024a); Viviers et al. (2023a); Schmidt et al. (2023); Long et al. (2021a); Zepf et al. (2023a; 2024); Hu et al. (2022); Viviers et al. (2023c); Savadikar et al. (2021); Philps et al. (2024)
HVAE Zhang et al. (2022c); Rafael-Palou et al. (2021); Kohl et al. (2019a); Baumgartner et al. (2019); Gantenbein et al. (2020)
@@ 6.2 Active Learning
The field of Active Learning aims to reduce the costly annotation procedure by careful selection of the most informative samples for training (Settles, 2009; Ren et al., 2021). Instead of randomly labeling data, the model directs expert time toward the most challenging examples, leading to more efficient learning. Active Learning is an extensively researched topic in both traditional (Cohn et al., 1996) and deep Machine Learning (Gal et al., 2017b), including uncertainty-based approaches applied to image-segmentation pairs. In such methods, it is assumed that predictions with high uncertainty provide the most informative samples for training. Moreover, this core idea extends to active perception, where an autonomous agent, like a robot or a drone, physically acts to resolve its own environmental uncertainty. For instance, instead of guessing what an obscured object is, it might move its camera to get a better view. In both cases, the system actively seeks the most valuable information, either from a human annotator or by interacting with the world, to improve its understanding and performance.
@@
Selecting the most informative samples for training contains two elements. Firstly, the uncertainty needs to be quantified. This is usually the predictive entropy (Kasarla et al., 2019; Shen et al., 2021; Xie et al., 2022) or variance (Ozdemir et al., 2021), but in some instances the uncertainty is estimated through Bayesian Active Learning by Disagreement (BALD) (Houlsby et al., 2011; Ma et al., 2024b; Shen et al., 2021). Simply stated, the reduction in posterior entropy indicates the informativeness of a data sample. In turn, this can also be formulated by the reduction in predictive entropy in the dual formulation. We can formally denote this as
@@
Note the similarity between the mutual information $I$ obtained from the predictive entropy decomposition in Equation (3). Secondly, to evaluate whether the active learning strategy is beneficial, the performance is mostly measured subject to increasingly stringent budget requirements (i.e. a percentage of the original dataset) (Yang et al., 2017; Sourati et al., 2018; Kasarla et al., 2019; Mahapatra et al., 2018; Sinha et al., 2019; Siddiqui et al., 2020; Li and Alstrøm, 2020; Kim et al., 2021; Shen et al., 2021; Burmeister et al., 2022; Ma et al., 2024b). Kahl et al. (2024) argue that Active Learning should be evaluated on image-level as humans do not annotate on pixel-level. Assuming a saturated performance on in-distribution data (training cycle $t_{1}$), the author propose to measure the improved relative performance, $C$, subject to a second training sample ($t_{2}$) from selected samples
@@
Table 2: Methods used for the application of Active Learning.
@@
MC dropout Kahl et al. (2024); Gorriz et al. (2017); Ozdemir et al. (2021); Hiasa et al. (2019); Li and Alstrøm (2020); Shen et al. (2021); Ma et al. (2024b); Gaillochet et al. (2023); Siddiqui et al. (2020); Sadafi et al. (2019)
@@
As can be seen in Table 2, most of the literature use MC dropout or ensembles. Furthermore, using the plain SoftMax activation has also been commonly used. Often, solely relying on the most uncertain samples can result in samples with limited diversity and therefore samples selected on the representativeness (Ozdemir et al., 2021; Shen et al., 2021; Wu et al., 2021; Mahapatra et al., 2018; Sinha et al., 2019). Another crucial modeling element pertains the aggregation of the pixel-level uncertainties. Therefore, closeness to boundary (Gorriz et al., 2017; Kasarla et al., 2019; Ma et al., 2024b) or specific regions (Gaillochet et al., 2023; Wu et al., 2021; Kasarla et al., 2019) are additionally incorporated when combining the uncertainty values across pixels. Siddiqui et al. (2020) determine the entropy through estimating the entropy across viewpoints in multi-view datasets. Gaillochet et al. (2023) argue the superiority of selecting samples based on batch-level uncertainty.
@@ 6.3 Model introspection
The provided uncertainty can be used to gauge prediction quality of a model. For example, an agricultural drone’s segmentation model provides a clear example. When distinguishing "crop" vs. "weed", it will segment healthy fields with low uncertainty, signaling a reliable output. However, for an anomalous patch of stressed plants, it will produce a segmentation with high uncertainty. This high score directly flags the prediction as low-quality and likely erroneous, preventing a costly automated action like misapplying herbicide. As evident, the described problem can also be considered as Out-of-Distribution (OOD) detection (Lambert et al., 2022; Holder and Shafique, 2021).
Table 3: Methods used for the application of Model Introspection.
@@
TTA Kahl et al. (2024); Wang et al. (2019b); Whitbread and Jenkinson (2022); Roshanzamir et al. (2023); Dahal et al. (2020)
Ensembling Jungo et al. (2020); Mehrtash et al. (2020); Czolbe et al. (2021); Holder and Shafique (2021); Jungo and Reyes (2019); Ng et al. (2022); Hann et al. (2021); Jungo and Reyes (2019); Linmans et al. (2020); Pavlitskaya et al. (2020); Valada et al. (2017)
MC Dropout Jungo et al. (2020); Kahl et al. (2024); Whitbread and Jenkinson (2022); Kendall et al. (2016); Kampffmeyer et al. (2016); Dechesne et al. (2021); Morrison et al. (2019); Qi et al. (2023); Eaton-Rosen et al. (2018); Jungo et al. (2018a; b); Roy et al. (2018a; b); Mehrtash et al. (2020); Nair et al. (2020); Sander et al. (2019); Hasan and Linte (2022b); Camarasa et al. (2021); Roshanzamir et al. (2023); Sedai et al. (2018); Seeböck et al. (2019); DeVries and Taylor (2018); Czolbe et al. (2021); Hoebel et al. (2019); Bhat et al. (2021); Dahal et al. (2020); Antico et al. (2020); Lambert et al. (2022); Hasan and Linte (2022a); Jungo and Reyes (2019); Ng et al. (2022); Iwamoto et al. (2021); Hoebel et al. (2020)
@@
VAE Chotzoglou and Kainz (2019); Viviers et al. (2023b); Czolbe et al. (2021); Bian et al. (2020)
@@
The relationship between uncertainty and model accuracy has been formalized by Mukhoti and Gal (2018) through the Patch Accuracy vs Patch Uncertainty (PAvPU) metric. Firstly, the accuracy given a certain prediction, $p(\text{A}|\text{C})$, and secondly, the uncertainty given an inaccurate prediction $p(\text{U}|\text{I})$. Given a threshold $u_{T}$ that distinguishes certain from uncertain pixels or patches, we can define pixels that are accurate and certain, accurate and uncertain, inaccurate and certain, inaccurate and uncertain, denoted by $u_{\text{ac}}$, $u_{\text{au}}$, $u_{\text{ic}}$, $u_{\text{iu}}$, respectively. Consequently, the authors combine $p(\text{A}|\text{C})=n_{\text{ac}}/(n_{\text{ac}}+n_{\text{au}})$ and $p(\text{U}|\text{I})=n_{\text{ic}}/(n_{\text{ic}}+n_{\text{iu}})$ to obtain the Patch Accuracy vs Patch Uncertainty (PAvPU) metric, defined as
@@
It can be noted that uncertainty is usually only obtained on pixel basis, while crucial information can be present in structural statistics. Hence, Roy et al. (2018b) propose to use the Coefficient of Variation (CoV), which addresses this by measuring structural uncertainty through dividing the volume variance over the mean for all samples. The authors also evaluate structural uncertainties by assuming predictions with the pair-wise average overlap between all respective samples.
@@
Considering Table 3, it can be observed that parameter modeling methods such as MC Dropout and Ensembles are by far the most popular choice of uncertainty encapsulation method. However, Concrete dropout (Rumberger et al., 2020b; Mukhoti and Gal, 2018), M-heads (auxiliary networks) (Jungo and Reyes, 2019; Linmans et al., 2020), MoE (Pavlitskaya et al., 2020) Laplace Approximation (Zepf et al., 2023b), Variational Inference in 3D (LaBonte et al., 2019), and test-time augmentation (Dahal et al., 2020) have also been used for this application. Moreover, including information on localized uncertainty to the training objective has shown to improve generalization capabilities (Ozdemir et al., 2017; Bian et al., 2020; Iwamoto et al., 2021; Li et al., 2021).
@@ 6.4 Model generalization
As mentioned earlier in Section 5.1, sampling new parameter permutations often improves segmentation performance, due to the model combining effect. In contrast to other downstream tasks, evaluation does not require specialized metrics besides quantifying the relative change in model performance. For instance, dropout layers at the deeper decomposition levels of the SegNet (Badrinarayanan et al., 2015) has shown to improve model performance (Kendall et al., 2016). MC Dropout, and specifically Concrete Dropout (Mukhoti and Gal, 2018; Rumberger et al., 2020b)), has also led to improved performance. The improved generalization from ensembling has also shown to produce more calibrated outputs (Mehrtash et al., 2020). Ng et al. (2022) found that ensembling results in the best performance improvement, while BBB is more robust to noise distortions. In other work, orthogonality within and across convolutional filters of the ensemble is enforced through minimizing their cosine similarity, which reaped similar merits (Larrazabal et al., 2021a). Ensembling has also been performed using M-Heads (Hu et al., 2023; Jungo and Reyes, 2019). Nonetheless, individual models in conventional ensembles receive data in an unstructured manner. Hence, specific subsets of the data can also be assigned to particular models with MoEs (Pavlitskaya et al., 2020; Ji et al., 2021; Gao et al., 2022). Generative models, though typically used for other applications, have also led to performance improvements across a wide range of studies (Saha et al., 2021a; Wolleb et al., 2021; Wu et al., 2023b; Amit et al., 2021; Bogensperger et al., 2023; Zbinden et al., 2023; Viviers et al., 2023b; Zepf et al., 2023a; Hu et al., 2022).
Key takeaways: Parameter distribution methods are commonly utilized for Active Learning, Model Introspection, and Model Generalization, while feature distribution modeling is employed for Observer Variability. This observation suggests the irreducibility of uncertainty tied to parameter-level modeling, while reducibility aligns with feature-level modeling.
@@
MC dropout Mukhoti and Gal (2018); Rumberger et al. (2020b); Saha et al. (2021a); Zhang et al. (2022a); Whitbread and Jenkinson (2022); Huang et al. (2018); Mukhoti et al. (2020); Ng et al. (2022); Wickstrøm et al. (2020)
Ensembling Ng et al. (2022); Larrazabal et al. (2021a); Kamnitsas et al. (2018); Saha et al. (2021b); Hu et al. (2023); Ji et al. (2021)
TTA Wang et al. (2019a); Rakic et al. (2024); Whitbread and Jenkinson (2022)